3x + y = 4
x + y = 0
Determine the solution to the system and check your solution in both equations. Input the solution to the system, putting the solution in ordered pair (x, y) form.
1
Expert's answer
2012-10-19T09:07:55-0400
from the second equation x=-y then we use it in the first one
3x+y=4 so
3*(-y)+y=4 -2y=4 y=-2 then x=2

check this solution 3*2+(-2)=4 and
2+(-2)=0
Answer: (2,-2)